OpenStudy (anonymous):

Factor completely p4 - 2p2 + 1. (p2 - 1)2 (p - 1)(p + 1)(p2 + 1) (p - 1)(p + 1)(p2 - 1) (p - 1)2(p + 1)2

OpenStudy (anonymous):

if you have the multiple choice, why don't you just work each answer and figure out which one it is. another hint. (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

thnks.

OpenStudy (anonymous):

also, (a^2-b)=(a+b)(a-b)

OpenStudy (hoblos):

(p2 - 1)2

OpenStudy (anonymous):

a^2-b^2 * my bad

OpenStudy (anonymous):

if you need the answer, its the last one

OpenStudy (hoblos):

it is the first

OpenStudy (anonymous):

the second one is the only one that doesnt give you the original equation. the other 3 give the answer, but the last one is fully factored.

OpenStudy (anonymous):

trust me. it is the last one if you want fully factored.

OpenStudy (hoblos):

yup tutordiffeq is right :D sorry